OpenStudy (anonymous):

Iron has a density of 7.86 g/cm3 (1 cm3=1 mL). Calculate the volume (in dL) of a piece of iron having a mass of 4.54kg . Note that the density is provided in different units of volume and mass than the desired units of volume (dL) and the given units of mass (kg). You will need to express the density in kg/dL before calculating the volume for the piece of iron.

OpenStudy (paxpolaris):

\[4.54kg \times {1cm^3 \over 7.86 g}\times{ 1mL \over 1cm^3} \times {1 dL \over 100 mL} \times {1000 g \over 1 kg}\]